### 10 Dec 2023 v3

# Pointer Networks Trained Better via Evolutionary Algorithms

Abstract
Pointer Network (PtrNet) is a specific neural network for solving Combinatorial Optimization Problems (COPs). While PtrNets offer real-time feed-forward inference for complex COPs instances, its quality of the results tends to be less satisfactory. One possible reason is that such issue suffers from the lack of global search ability of the gradient descent, which is frequently employed in traditional PtrNet training methods including both supervised learning and reinforcement learning. To improve the performance of PtrNet, this paper delves deeply into the advantages of training PtrNet with Evolutionary Algorithms (EAs), which have been widely acknowledged for not easily getting trapped by local optima. Extensive empirical studies based on the Travelling Salesman Problem (TSP) have been conducted. Results demonstrate that PtrNet trained with EA can consistently perform much better inference results than eight state-of-the-art methods on various problem scales. Compared with gradient descent based PtrNet training methods, EA achieves up to 30.21% improvement in quality of the solution with the same computational time. With this advantage, this paper is able to at the first time report the results of solving 1000-dimensional TSPs by training a PtrNet on the same dimensionality, which strongly suggests that scaling up the training instances is in need to improve the performance of PtrNet on solving higher-dimensional COPs.
1 Introduction
The Pointer Network (PtrNet) [1] is designed to tackle NP-hard Combinatorial Optimization Problems (COPs). PtrNet follows the typical machine learning paradigm: it is trained offline on a set of COP instances and then applied online to infer solutions for new instances from the same problem class. Due to its feed-forward inference procedure and the availability of hardware acceleration techniques [2], PtrNet can solve incoming instances in seconds or less[3]. This makes PtrNet superior to traditional mathematical programming[4] and meta-heuristic search methods [5] in many scenarios with real-time demands, such as Energy Management [6], Unmanned Aerial Vehicle System [7], and Tasks Scheduling in Cloud Computing [8].
For this advantage, PtrNet has drawn increasing attention from the research community recently. Its promising performance has been verified in some classical COPs like Travelling Salesman Problems (TSPs) [9], Vehicle Routing Problems (VRPs)[10], Job Schedule Problems (JSPs)[11] and Max Clique Problems (MCPs) [12]. For example, when solving TSP instances with 50 nodes , the Hybrid PtrNet reduces the reference time from 1120s to 4s and obtains better results compared with OR-Tools [3]; when solving VRP instances with 30 customs and 6 charging stations, the PtrNet-based methods hierarchical GPN could achieve similar performance by consuming one-tenth less computing time of the state-of-the-art (SOTA) methods [10].
On the other hand, one major drawback of PtrNet is the lower quality of the results compared with traditional methods, especially on larger-scale COPs, as systematically evaluated by [13] . For example, the quality gap of result between PtrNet and the well-established solver Lin-Kernighan heuristic (LKH) [14] degenerate significantly from 19.81% to 41.44% on TSP with 250 nodes and 1000 nodes [15], respectively. Although recent works have proposed more advanced models like Attention [16] and POMO [17] , they still cannot resolve the above issue. On this basis, this paper focuses on the training method of the neural networks (NNs) for COPs, rather than the model design. We initially opted for the PtrNet over the more complex models for simplicity and generality. The straightforward architecture of PtrNet provides a clear foundation for implementing and understanding the effects of EAs in training NNs, and establishes a baseline before exploring more intricate models.
Problem Method Research Training Testing
Dimensionality Dimensionality
TSP PtrNet+SL [1] 50 50
PtrNet+RL [18] 100 100
Graph PtrNet+RL [15] 50 1000
PtrNet+REINFORCE [19] 100 100
PtrNet+SL [20] 150 150
Hybrid PtrNets [3] 100 100
VRP PtrNet+RL [21] 100 100
PNSP-VNS [9] 100 130
PtrNet+RL [11] 100 100
PtrNet+REINFORCE [19] 100 100
JSP PtrNet+RL [11] 50 50
MCP PtrNet+SL [12] 45 50
Table 1: Training and testing dimensionalities of existing PtrNet-based methods.
Existing training methods for PtrNet can be roughly classified into two categories based on how the training errors are obtained, i.e., Supervised Learning (SL) and Reinforcement Learning (RL). The major difference of these training methods lies in the origination of their error signals for optimizing the network parameters. SL adjusts the parameters by comparing the inferred solutions with the ground truth labels, i.e., the qualities of the best known solutions [1], thereby making the quality and acquisition of these labels a bottleneck for such methods. In contrast, for RL, the error signal is derived from the rewards or penalties by intermediately interacting with the environment, i.e., the evaluation function of solution qualities [18]. Despite of the error signals, both methods fundamentally relies on the gradient descent to optimize network parameters. As revealed by [22], unfortunately, a major challenge of gradient descent based methods is that their search processes are probe to be trapped by local optima. In other words, the sub-optimal results of PtrNet could possibly stem from the issue that gradient methods cannot locate a sufficiently favorable optimum within the parameter space of PtrNet. In recent years, well accepted studies have examined the potential superiority of Evolutionary Algorithms (EAs) over gradient methods for optimizing the parameters of Deep Neural Networks (DNNs) for specific problems, e.g., Deep RL for video games [23] and Neural Architecture Search [24]. A pertinent question arises that "could EA also benefit the training of PtrNets?". This paper aims to, at the first time, provide a set of sophisticated empirical evidence to answer this question.
In the framework of training DNN with EAs, EA iteratively samples a population of individuals (i.e., networks) in the parameter space based on certain randomized search operators [25]. This sampling process is essentially driven by a simulated search gradient formed by multiple evaluated individuals [26]. As a result, EA is more likely to escape from the local optima compared to gradient descent methods, as the simulated search direction is less greedy. Moreover, the population-based search process can further provide the possibility of exploring multiple different regions of the parameter space with the aid of population diversity [27]. This parallel exploration ability can not only enhance the search efficiency in a way of divide-and-conquer, but also be accelerated by parallel computing [28].
In order to verify the aforementioned advantages of EAs, we have designed sophisticated computational experiments on training PtrNet for TSPs. The experiments specially focus on an important training challenge. That is, existing training methods are hardly to train PtrNet effectively on 1000-dimensional instances, which are more realistic to real-world problem scales [29]. Albeit PtrNet is able to handle varied dimensional instances on its recurrent network structure, the quality of solutions deteriorates as the difference between the testing dimensionality and training dimensionality increases, as empirically revealed by [30]. For instance, PtrNet trained on 150 dimensions, when used to solve a problem of 1000 dimensions, has a disparity of up to 41.44% compared to the solution obtained by LKH [15]. As suggested in these works, to solve 1000-dimensional TSPs, it is more natural and effective to train PtrNet on the same dimensionality. Nevertheless, Table 1 illustrates that current methods predominantly focus on various COPs around 150 dimensions or less. While a few methods attempt solving 1000-dimensional COPs, their training instances are still no more than 150 dimensions, and the results all faced the issue of low quality. This situation remains the same for recently proposed Attention and POMO models, which underscores the importance of training and solving at similar dimensions to enhance the solution quality for high-dimensional COPs. However, this will result in more local optima in the parameter space and lead to a higher risk that the search process is trapped [31]. On this basis, this challenge is a natural scenario for verifying the merits of EAs for training PtrNet.
We employed an EA called Negatively Correlated Search (NCS) [27] to instantiate the EA training framework, denoted as PtrNet-EA. PtrNet-EA is extensively compared with eight methods, including three distinct training methods for PtrNet and five SOTA non-PtrNet methods, comprising two learning-based methods and three heuristic methods. A series of empirical studies have been conducted on TSPs and the major results are as follows:
  • We assess the performance of learning-based algorithms on inferring 1000-dimensional TSP instances, trained with different dimensional instances within a fixed computational time budget. It shows that when the training dimensionality increase, the solution quality of the compared algorithms even degenerate by 10.57%, while PtrNet-EA enjoys 2.68% improvements and consistently outperforms the compared ones.
  • We test the computational time needed by all algorithms to reach the same targeted solution quality on 1000-dimensional TSP datasets. It is found that PtrNet-EA is able to consume much fewer computational time than the compared algorithms.
  • We study the benefit of population size to the performance. By enlarging the population size up to 90, PtrNet-EA gains 4.74% further improvement than PtrNets-EA with a population of 5.
  • We test how the diverse population of PtrNet-EA helps to solve TSPs, with making all the PtrNets in the final population as a portfolio to infer each incoming instance and recording the best result. Empirical results witness 1.99% further improvement of PtrNet-EA, and finally achieves in total 30.21% improvement over gradient-based PtrNet trainers, within the same acceptable computational time budget.
The contributions of this paper can be summarized as follows:
  • This is the first work empirically assesses the advantages of training PtrNets with EA over traditional gradient based training methods. This provides new insights to improve neural solvers with EAs.
  • This is the first work empirically verifies that training PtrNet on 1000-dimensional instances can significantly benefit the inference on 1000-dimensional instances. This suggests that scaling up training instances is in need to improve the neural solvers on large-scale problems.
2 Related Work
In this section, we review seminal works related to our study. We first discuss the application and evolution of PtrNets in COPs, highlighting their advancements and significance. Then, we delve into the application of EAs for optimizing NNs, emphasizing their potential advantages over gradient-based methods.
2.1 PtrNets for Combinatorial Optimization Problems
The earliest network for COPs was Hopfield network [32]. Since then, researchers have invested decades in this field to address the generalization and efficiency issues of neural combinatorial solvers. Recurrent Neural Networks (RNNs) [33] solves the problem that Hopfield Network needs to be retrained for each instance. Unfortunately, RNNs cannot handle COPs with varied dimensions, thus their generalization ability are still restricted. With the continuous development of sequence-to-sequence (S2S) paradigm, the PtrNet, trained with SL, to learn the mapping between the input of COPs instances as a feature sequence and the output as a solution sequence. Under the S2S paradigm, the varied dimensions issue of RNNs is solved, and PtrNet ushered in a new era of neural combinatorial solvers. Thereafter, [12] applied PtrNet to solve Max-Cut problems(MCPs), significantly reducing the required time and computational resource for solving it with heuristic methods, demonstrating the versatility and potential of PtrNet in addressing COPs. [9] proposed a PtrNet solution pool which employs a set of PtrNets to ensure that the breadth of the search does not disappear with the iterations. [20] introduced a heterogeneous PtrNet that adapts the encoder to the spatial-temporal feature of Traveling Officer Problems, demonstrating superior performance in capture rate and travelling distance on real-world datasets compared to existing heuristic and deep learning methods.
In addition to the aforementioned SL-based PtrNet, the PtrNet has been continuously improved under the framework of RL (PtrNet-RL) since its inception. [18] trained the PtrNet by the famous Actor-Critic algorithm [34]. They demonstrated that even trained with optimal labels, PtrNet-SL was rather poor compared to PtrNet-RL. Subsequently, [21] made improvements on this basis. With graph embedding to handle non-Euclidean data, [15] proposed the Graph Pointer Networks (GPN) trained by RL. [19] introduced REINFORCE with greedy rollout baseline to the PtrNet-RL framework. They employed a multi-head attention mechanism to capture the relationship between the nodes in COPs, thus the results are significantly improved for small-scaled TSP. [35] leveraged PtrNet for the multi-objective TSP, and the experimental results highlight its robust generalization, fast solving speed and competitive solution quality. [3] used a hybrid PtrNet with time features to solve TSP with time windows, and their model shows superior performance over existing methods with 20, 50, and 100 nodes.
2.2 EAs for Optimizing Neural Networks
Beyond the listed gradient-based methods above to train PtrNets, there is a substantial body of research discussing EAs as a highly competitive alternative. These population-based algorithms offer advantages over gradient-based methods by being less susceptible to local optima, operating without the need for gradient information, and aptly handling problems without a clear objective function. Given these advantages, EAs were introduced for training NNs as early as the 1990s [36]. In fact, the combination of EAs and NNs have achieved significant breakthroughs in various problems [37, 38]. [39] proposed an Evolution Strategies (ES) to train the RL agent, in which the agent updated by the best agent that sampled from a Gaussian distribution noise of last agent. Their method is fifteen times faster than the other gradient-based methods on MuJoCo locomotion tasks. [28] proposed Negatively Correlated Natural ES to directly training a Deep Convolution Network with 1.7 million parameters for playing Atari games, and achieved twice acceleration over compared gradient-based methods by incorporating co-evolution mechanisms [40].
Currently, there are some researchers explore to integrate EAs and neural solvers. [41] uses NN to initialize the population of a genetic algorithm and further evolves the population through the genetic algorithm to obtain the final policy. [42] and [43] solve large-scaled TSP and Multi-objective TSP with EAs and PtrNets, respectively. However, due to the large parameter space of PtrNets, there has been limited research so far on directly applying EAs on PtrNets. The most related work to this paper was proposed in [43], where a multi-objective evolutionary algorithm based on decomposition and dominance is presented to train the PtrNet on TSP and Knapsack problems. Their works mainly focus on how to handle multiple objectives of the COPs variants or decompose large-scaled COPs instances, rather than focus on the EAs as an effective alternative of RL in training PtrNets. In this work, we at the first time thoroughly discuss the advantages mentioned in the previous section.
3 Methodology
In this section, we begin by providing a concise introduction of the structure and trainable parameters of the PtrNet, and then introduce the PtrNet-EA framework. Lastly, we implement the PtrNet-EA framework by employing the NCS algorithm.
3.1 Parameter Representation of PtrNet for EA
PtrNet, a specific neural architecture, comprises two Long-short Term Memory (LSTM) modules, encoder and decoder. The encoder network sequentially processes the input sequences, which represents nodes, and coverts it into a series of latent memory states. The decoder network also maintains its latent memory state, and it employs a pointing mechanism to generate a distribution over the next chosen nodes. The parameters in both of the encoder and decoder LSTM modules are trainable, and the number of the them are listed in Table 2, where d denotes the dimension of the hidden layer, which is 256. The workflow of PtrNet is shown in Fig.1, and the detailed structure of the encoder and decoder of PtrNet are shown in Fig.2 and Fig.3.
Figure 1: Flowchart of PtrNet. The green arrows denote the output pointer and the red arrows denote the pointers disabled by the masking mechanism.
Figure 2: The structure of the encoder module in PtrNet.
Figure 3: The structure of the decoder module in PtrNet.
Components Number of parameters
Encoder $13107205layers\times 4\times(d\times d)=$ 1310720
Decoder $13107205layers\times 4\times(d\times d)=$ 1310720
Attention Mechanism $d\times d=262144$
Table 2: Trainable parameters in PtrNets
3.2 Framework of PtrNet-EA
An evolutionary framework for NNs is an optimization approach that draws inspiration from the natural process of evolution, aiming to optimize the structure and/or parameters of NNs through iterative generations. For neural solvers addressing COPs, their structure remains constant during training, with only the network weights requiring adjustment. Based on this, in this work the evolutionary framework is employed to optimize the parameters of the PtrNet while keeping the architecture fixed. The following workflow describes how EAs optimize the parameters, each individual in the population denotes the parameters of a PtrNet.
  1. 1.
    Initialization. Generate an initial population of candidate PtrNets, each with a distinct parameters. The population size is predefined, and the architectures is based on previous works.
  2. 2.
    Fitness Evaluation. Evaluate the performance of the individuals using a predefined objective function.
  3. 3.
    Reproduction. The search operators of EAs, e.g., selector, crossover, and mutation, are used for generating the next population.
  4. 4.
    Termination Condition. Repeat step 2 and 3 until a stopping criterion is met, e.g., a predefined time budget, reaching a performance threshold or detecting convergence.
Upon completion of the evolutionary process, the PtrNet with the highest fitness is considered the optimized policy for inferring COPs.
3.3 Concrete EA-based Training Algorithm
In this paper, we choose NCS to instantiate the PtrNet-EA. NCS is a type of EA with explicitly enhanced diversity and has been applied to solve numerous practical problems [28, 40]. NCS consists of multiple parallel search processes. Each search process is driven by a Gaussian mutation operator [44], which generates the new individual $x_{i}^{\prime}$ by sampling from $\mathcal{N}\left(x_{i},\sigma_{i}\right)$, where $\sigma_{i}$ is the standard deviation of the Gaussian distribution $\mathcal{N}$ and represents the search step-size. The individuals are expected to be with high quality and distant from each other to form a parallel exploration of multiple regions of the parameter space. In NCS, the distance is defined as Bhattacharyya distance [45] by Eq.(1), where $p_{i}$ represents the distributions corresponding to other individuals. NCS forces the population seeking different optimal areas by trading off between the distance and the objective fitness values with an adaptive parameter $\lambda_{t}$. The $\sigma_{i}$ is updated based on the 1/5 success rule [46]. That is, $\sigma_{i}$ will be increased to enhance the exploration if 20% of the offspring $x_{i}^{\prime}$ is better than its parent $x_{i}$ within $t_{e}$ iteration; otherwise, it will be reduced for favoring exploitation, according to Eq.(2), where $c_{i}$ counts the replaced individuals. The pseudo-code is given in Algorithm , and at the end of the algorithm, the final population obtains all the PtrNet individuals with good performance and diversity. For more details of NCS, please refer to [27].
$Corr\left(p_{i}\right)={\mathbf{min}}_{j}{\{D_{B}(p_{i},p_{j})|j\neq i\}}$ (1)
4 Empirical Studies
This section assesses the potential of PtrNet-EA in the following aspects:
  • Whether EAs can train the PtrNets better than traditional gradient-based methods within an acceptable training time budget?
  • How would the parallel computing accelerate EAs in training pointer network?
  • How does the population diversity facilitate the performance of pointer network?
4.1 Travelling Salesman Problem
In this paper, the TSP is selected as the benchmark problems as it is commonly tested as a typical NP-hard combinatorial optimization problems [1]. In general, given a set of nodes in a graph and their coordinates, the goal of TSP is to find the shortest route that visits each node exactly once and returns to the starting node.
Formally, let $X=\{X_{1},X_{2},\dots,X_{n}\}$ be $n$ nodes in the TSP instance. The goal is to optimize a permutation over the nodes that minimises the path length
$L(\pi,X)=\sum\nolimits_{i=1}^{n}\left\|X_{\pi(i)}-X_{\pi(i+1)}\right\|_{2}\$ (2)
subject to:
$\pi(1)=\pi(n+1),$
$\pi(i)\neq\pi(j),$
$\pi(i)\in\{1,2,\dots,n\}$
Where $\sigma\left(i\right)$ is the index of node $i$ in the path sequence. Each node $X_{i}$ is represented by the coordinates where it is located. That is $X_{i}=(x_{i},x_{2},\ldots,x_{d})$, where $d$ is the dimension of the coordinates. In this paper, two-dimensional Euclidean coordinates are used for the representation, or $d=2$.
For the TSP instances, existing real-world benchmarks have been found insufficient for training PtrNets in terms of number of instances [15]. Therefore, following the method by [15], we synthesize three sets of TSP instances with 100 nodes, 500 nodes, and 1000 nodes, denoted as TSP100, TSP500, and TSP1000, respectively. Each set consists of 1 million training instances and 2000 testing instances, which are sampled from the same instance distribution by randomly distributing the x, y coordinates of the nodes in the real-valued interval of $\left[0,1\right]^{2}$.
4.2 Algorithms Settings
Suggested by [47], three types of well-established combinatorial solvers are selected as the compared algorithms. The first type contains the PtrNet-based methods, including the original PtrNet trained by SL (PtrNet-SL), the Graph Pointer Network with 2-opt (GPN-RL) [15], the pointer network trained by RL (PtrNet-RL) [18]. The second type includes two non-PtrNet learning-based solvers, i.e., the Structure to Vector Deep Q-learning (S2V-DQN) [48] and Variable Strategy Reinforced LKH (VSR-LKH) [49]. S2V-DQN is a deep learning architecture over graphs, and VSR-LKH is a hybrid of RL and LKH. The third type consists of the heuristic search based SOTA methods, i.e., the 2-opt [50], the Google OR-Tools with saving algorithm, and Focused Ant Colony Optimization (FACO)[51].
By comparing with the first two types, it aims to show the learning ability of the proposed PtrNet-EA. By comparing with the third type of methods, it emphasizes the superiority of PtrNet-EA over the traditional solvers within an acceptable time budget. The parameters and implementations of the above compared algorithms exactly follow the source codes released by [15]. The hyperparameters of PtrNet-EA are listed in Table 3. The embedding size, weight size, and batch size are set to the same with [1]. There are approximately 2.9M weights to optimize. The population size is set to 5 while comparing with SOTAs, and later set to 50 and 90 for parallel assessment.
Component Hyperparameter Value
PtrNet Embedding Size 32
Hidden State Size 256
Layer Number 5
Batch Size 256
NCS Max Time Steps 8000
Update frequency 10
Population Size [5, 50, 90]
Table 3: Hyperparameters used in PtrNet-EA.
4.3 Experimental Settings
This work aims to study how to train PtrNets with 1000-dimensionality TSP efficiently. We set the time budget to 500 minutes for all the algorithms for fairness. Indeed, any other time budgets larger than 500 minutes do not mean unacceptable. We simply approximate it around one working day of 8 hours for the consideration of real-world activities. More specifically, for learning-based solvers, the training phase is set to 500 minutes, as the testing phase is usually very efficient and thus the testing time can be omitted. For the heuristic search-based methods, we allow them to solve the instances in 500 minutes. Unless specified, all experimental results are the average of three independent runs. All experiments are conducted on a workstation with 320GB RAM and 96 cores (4.0GHz, 71.5MB L3 Cache), running Ubuntu 20.04. The experiments are divided into four Groups and will be described in detail.
Group I. We compare the proposed PtrNet-EA with the 8 chosen algorithms. All the learning-based solvers are first trained with 1 million training instances of TSP100, and then tested on 2000 testing instances of TSP100, TSP500, and TSP1000, respectively. The 2-opt, OR-Tools and FACO are directly used to solve the 2000 testing instances, as they do not involve the training step. The output path lengths of 2000 testing instances are averaged as the final performance of each algorithm.
Group II. A question naturally arises why existing works do not train the networks at the same scale with the testing instances. To clarify this, we train all the learning-based solvers on the training instances of TSP1000 and test them on the testing instances of TSP1000. In order to further analyze the superiority of PtrNet-EA on TSP1000, we use the testing performance of PtrNet-EA trained in the above-mentioned 500 minutes time budget as the baseline and run the compared algorithms with a 10x time budget. For every 50 minutes, we record the performance of compared algorithms on the testing instances. If the performance of a compared algorithm reaches the baseline, we record the consumed running time for that algorithm. Otherwise, we record the final performance of that algorithm stopped at 5000 minutes. This group of comparisons verifies whether EAs could help pointer networks achieve better performance in terms of running time on larger problems.
Group III. To verify how the parallel computing accelerates EAs in training PtrNets, the population size N of PtrNet-EA is set to 5, 50 and 90, respectively. For each of them, N processors are allocated to parallelize the training of individual policies, and the training time budget is also set to 500 minutes. By this means, it is expected to see that: 1) all the three algorithms will run for the same iterations; 2) due to the larger population sampled in each iteration, the PtrNet-EA with larger population size will achieve better results. If this empirically holds, it is promising to say that by utilizing more parallel computing resources, EAs can train the PtrNets even better within the same time budget.
Group IV. By treating the final population of evolved PtrNets as a policy portfolio, the complementary strengths of the individual policies can be leveraged to further improve the performance on COP instances. We verify this by comparing the best policy with policy portfolio, denoted as PtrNet-EA-P, on TSP1000.
4.4 Results and Analysis
Group I. When compared with PtrNet-SL, PtrNet-EA has 18.35%, 18.70% and 21.61% improvements on TSP100, TSP500 and TSP1000, respectively. When compared with RL-based PtrNets, PtrNet-EA has at least 11.95%, 17.23% and 10.91% improvements, respectively. These comparisons show that the PtrNet-EA has a great learning ability. When compared with other SOTA heuristic methods, the results have been improved by 4.40%, 4.27% and 2.22%, , which means that PtrNet-EA is able to achieve better solution in time-constrained tasks. The detailed results are shown in Table 4.
Methods TSP100 TSP500 TSP1000
PtrNet-SL $10.29\pm 0.60$ $22.61\pm 0.85$ $33.89\pm 0.36$
PtrNet-RL $10.60\pm 0.98$ $22.21\pm 0.92$ $31.44\pm 0.53$
GPN-RL $9.54\pm 0.45$ $22.40\pm 0.24$ $29.82\pm 0.27$
S2V-DQN $11.09\pm 1.12$ $22.21\pm 0.30$ $27.66\pm 0.36$
VSR-LKH $8.94\pm 0.19$ $21.22\pm 0.56$ $27.30\pm 0.43$
OR-Tools $9.41\pm 0.09$ $21.26\pm 0.82$ $29.50\pm 0.85$
2-opt $9.02\pm 0.67$ $20.16\pm 0.26$ $29.21\pm 0.41$
FACO $8.79\pm 0.16$ $19.20\pm 0.21$ $27.17\pm 0.52$
PtrNet-EA 8.40 $\pm$ 0.13 18.38 $\pm$ 0.57 26.57 $\pm$ 0.56
Table 4: Comparison of methods trained on TSP100 and Tested on TSP100, TSP500 and TSP1000, by $Avg\pm Std$.
Next, we compared all the learning-based methods trained on TSP1000 and tested on TSP1000. Not surprisingly, PtrNet-EA still maintains the leading position in this comparison. As Table 5 shows, some of the other methods have performance loss in this experiment especially the RL-based ones. However, PtrNet-EA benefits from the high-dimensional datasets, and the quality of the solution is improved by 1.74% in comparison with training on TSP100.
Methods Trained on TSP100 Trained on TSP1000
PtrNet-SL 33.89 $\pm$ 0.36 32.75 $\pm$ 0.72
PtrNet-RL 31.44 $\pm$ 0.53 31.98 $\pm$ 0.80
GPN-RL 29.82 $\pm$ 0.27 29.29 $\pm$ 0.81
S2V-DQN 27.66 $\pm$ 0.36 30.62 $\pm$ 0.70
VSR-LKH 27.30 $\pm$ 0.43 32.35 $\pm$ 0.20
PtrNet-EA 26.57 $\pm$ 0.56 25.86 $\pm$ 0.58
Table 5: The solution quality of learning-based models trained on TSP100 and TSP1000, and Tested on TSP1000, by $Avg\pm Std$.
Group II. The results are shown in Fig.4. It is obvious that the compared methods need over 10x the time budget to achieve the same performance of PtrNet-EA, and fully illustrates the efficiency of EA in training PtrNets.
Figure 4: 10x time budget for compared methods training on TSP1000 to reach the PtrNet-EA’s performance in 1x time budget.
Group III. Based on the parallel nature of EAs, we believe that a larger population can achieve better search results. Fig.5 shows the effect of population size to PtrNet-EA. From the convergence curve, we can clearly figure out that larger population size could reach the specified performance faster, and within a given time, larger one performs better, which confirmed our conjecture. By removing the diversity control element from PtrNet-NCS(pop5), the PtrNet-EA(noDiv5) method got a huge drop of 26.8% (see Table 6), which also verified the importance of population diversity in PtrNet-EA indirectly. It is worth mentioning that in parallel condition, there is almost no extra training time in PtrNet-EA.
Figure 5: Convergence curve of PtrNet-EA with different population size and PtrNet-RL.
Figure 6: Diversity of the policy portfolio (pop5).
Method PtrNet-EA PtrNet-EA-P
PtrNet-EA(noDiv5) 32.12 31.87
PtrNet-EA(pop5) 25.33 25.12
PtrNet-EA(pop50) 24.49 24.21
PtrNet-EA(pop90) 24.13 23.65
Table 6: Performance difference between PtrNet-EA and PtrNet-EA-P.
Group IV. Last but not least, the policy portfolio can also make some contributions. When we use all policies in the last population as a portfolio, each policy has the potential to obtain the optimal solution on specific problem instances. As shown in Fig.6, 34% best solutions of the 2000 testing instances are obtained by the policies in the portfolio other than the best policy. As much as 1.99% improvement is introduced by the policy portfolio (see Table 6), and the overall improvement reached an astonishing 30.21% compared with the original PtrNet-SL.
5 Conclusion
In this paper, we delved deeply into the performance constrains observed in the PtrNets when solving COPs. Recognizing the limitations of gradient descent methods, primarily their susceptibility to being trapped in local optima, our study postulated the potential effectiveness of EAs as a competitive alternative. To verify this, we systematically validated these through sophisticated four groups of empirical experiments, and the results consistently demonstrated the superior capability of EAs in training PtrNets.
Therefore, our work provides a new insight to both fields of EAs and PtrNets, which indicating that EAs could train PtrNets more proficiently than gradient-based methods, e.g., SL and RL. This suggests that for applications that leverage PtrNet in solving large-scale COPs, utilizing EAs for training should be seriously considered.
Moreover, while our current research focused on the PtrNet, given the sequential structures also shared by the models like Attention and POMO, it is reasonable to induce that they can also benefit from similar evolutionary training framework. In the future, how to train models better with EAs will be studied as they appear more complex.